The 2021 income statement for Anderson TV and Appliance reported net sales of $360,000 and net income of $85,000. Average total assets for 2021 was $800,000. Shareholders’ equity at the beginning of the year was $500,000 and $40,000 was paid to shareholders as dividends. There were no other shareholders’ equity transactions that occurred during the year.

Show the DuPont framework's calculation of the three components of the 2021 return on equity for Anderson TV and Appliance.

Profit Margin (%)
Choose Numerator: / Choose Denominator: = Profit Margin
Net income / Net sales = Profit Margin
? / ? = ?
Asset Turnover
Choose Numerator: / Choose Denominator: = Return on Assets
Net sales / Average total assets = Asset turnover
? / ? = ? times
Equity Multiplier
Choose Numerator: / Choose Denominator: = Return on Equity
Average total assets / Average shareholders’ equity = Equity multiplier
$800,000 / ? = ? times
Return on equity ?
Answer #1

Show the DuPont framework's calculation of the three components of the 2021 return on equity for Anderson TV and Appliance.

Profit Margin (%)
Choose Numerator: / Choose Denominator: = Profit Margin
Net income / Net sales = Profit Margin
85000 / 360000 = 23.61%
Asset Turnover
Choose Numerator: / Choose Denominator: = Return on Assets
Net sales / Average total assets = Asset turnover
360000 / 800000 = 0.45 times
Equity Multiplier
Choose Numerator: / Choose Denominator: = Return on Equity
Average total assets / Average shareholders’ equity = Equity multiplier
$800,000 / 522500 = 1.53 times
Return on equity 23.61*1.53*.45 = 16.26%

Ending stockholder's equity = 500000+85000-40000 = 545000

Average equity = (500000+545000/2) = 522500